OpenStudy (anonymous):

Identify the correct slope and y intercept of the equation 6x - 3y = 12.

OpenStudy (countrygirl16):

y = mx + b, m being slope and b being the y intercept. so we want to find m. 6x - 3y = 12 divide by three 2x - y = 4 move 2x over -y = -2x + 4 multiply by -1 y = 2x - 4 slope (m) is 2